Find the slope of the line passing through the given points

(-4,-1),(-1,-7)

Answer by mananth(16949)

x1 y1 x2 y2
-4 -1 -1 -7

slope m = (y2-y1)/(x2-x1)
( -7 - -1 )/( -1 - -4 )
( -6 / 3 )
m= -2
The required line passes through both the points. So we select one point and plug it with the slope
Plug value of the slope and point ( -4 , -1 ) in
Y = m x + b
-1.00 = -2 * -4 + b
-1.00 = 8 + b
b= -1 + -8
b= -9
So the equation will be
Y = -2 x -9
m.ananth@hotmail.ca